# Hybrid Querying Over Relational Databases and Large Language Models

### v1

@@ Abstract.
Database queries traditionally operate under the closed-world assumption, providing no answers to questions that require information beyond the data stored in the database. Hybrid querying using SQL offers an alternative by integrating relational databases with large language models (LLMs) to answer beyond-database questions. In this paper, we present the first cross-domain benchmark, SWAN, containing 120 beyond-database questions over four real-world databases. To leverage state-of-the-art language models in addressing these complex questions in SWAN, we present, HQDL, a preliminary solution for hybrid querying, and also discuss potential future directions. Our evaluation demonstrates that HQDL using GPT-4 Turbo with few-shot prompts, achieves 40.0% in execution accuracy and 48.2% in data factuality. These results highlights both the potential and challenges for hybrid querying. We believe that our work will inspire further research in creating more efficient and accurate data systems that seamlessly integrate relational databases and large language models to address beyond-database questions.
@@ 1. Introduction
The Relational model and SQL have achieve widespread acceptance and usage in data management systems. In particular, SQL continues to evolve by adding new syntax and features, enhancing its capabilities to meet the growing demands of modern data systems (Stonebraker and Pavlo, 2024). It is well known that database queries are evaluated under a closed domain assumption, meaning that the queries address aspects of the real world based solely on the data stored in the relational database management system (Reiter, 1988). While the direct approach is to say NO to beyond-database questions, researchers in our community have also investigated alternatives such as providing answers based on incomplete data with heuristic algorithms (Levy, 1996) and crowdsourcing (Franklin et al., 2011). In this paper, we explore the use of large language models to address these questions.
In the natural language processing community, open-domain question answering has been extensively studied, often encompassing a broader range of inquiries. This long-standing task aims to provide factual answers to natural language questions by drawing from large, unstructured collections of texts and documents, such as Wikipedia. By pre-training on large corpus of knowledge, large language models (LLMs) have demonstrated significant potential in providing world knowledge and performing complex reasoning (Hendrycks et al., 2020; Zhao et al., 2023). In this paper, we are specifically interested in beyond-database questions which have partial information grounded in the relation database. Unlike open-domain questions, beyond-database questions require integrating and reasoning with structured data from relational databases as well as drawing from large, unstructured collections of texts.
@@
A motivating example: Consider a simple database with a single table containing superhero information, as shown in Figure 1. The schema for the database is: superhero(hero_name, full_name). Suppose the user wants to list all the hero names from the Marvel Universe within the database. This would be considered as a beyond-database question, since the database contains relevant or partial information to the request but cannot directly provide the answer (i.e., which heroes are Marvel universe). On the other hand, large language models such as ChatGPT can be used to identify the publisher for each superhero characters. Assume we treat LLMs as a table containing the hero_name and the publisher. Then, a SQL query like: ”SELECT hero_name, full_name FROM llm JOIN superhero ON llm.hero_name = superhero.hero_name WHERE llm.publisher = ’Marvel’;” Hence, hybrid querying integrating relational databases and large language models offers a powerful approach for addressing beyond-database questions.
In this work, we propose SWAN, Solving beyond-database queries With generative AI aNd relational databases, the first hybrid query benchmark. SWAN comprises 120 beyond-database questions and spans four big databases, covering four diverse domains. The original databases and questions are from the recent Bird benchmark(Li et al., 2024), which is a benchmark for evaluating natural language to SQL translation. The data in these databases are collected from open-source relational databases in platforms such as Kaggle. Also, the proposed SWAN benchmark challenges large language models to both select values from a given list (e.g., choose publisher name from a list of predefined publishers) and generate free-form outputs (e.g., determine the city based on a street address). In addition to proposing the SWAN benchmark, we also introduce HQDL, a preliminary solution for answering beyond-database questions. We evaluate HQDL over the state-of-the-art large language models such as ChatGPT (gpt-3.5-turbo) (Ouyang et al., 2022) and the GTP-4 turbo (Open, 2023; Achiam et al., 2023). Our experimental results indicate significant challenges for current state-of-the-art models in generating factual data and accurately answer beyond-database questions with hybrid queries. With in-context learning (ICL), GPT-4 Turbo achieves 40.0% execution accuracy and 48.2% of its generated data is factually correct. Proposing a novel benchmark for beyond-database queries will focus attention and drive more research in this critical area. It will also encourage researchers to develop innovative solutions that are founded on solid well understood foundations. The evaluation results based on HQDL further highlight the need for more advancements in improving the accuracy and reliability of answering beyond-database questions with hybrid queries.
@@ 2. Background and Problem Statement
In this section, we provide a discussion on the preliminaries of large language models (LLMs) and introduce related work from both the database and the natural language communities for answering beyond-database questions.
2.1. Preliminaries on Large Language Models
@@
Large language models are frequently used to retrieve domain-specific knowledge, simplifying the information retrieval process by providing direct natural language answers. However, these models may exhibit hallucinations and factuality errors. Consequently, recently many researchers have focused on enhancing the capability of LLMs to provide factual information (Wang et al., 2023a). In-context learning (ICL) is a widely adopted strategy to enhance the factual accuracy of generated content. ICL enables a language model to learn from example demonstrations within its context to improve performance (Brown et al., 2020).
2.2. Hybrid Query Task Formulation
SELECT T1.full_name, T1.superhero_name
FROM superhero AS T1 INNER JOIN llm AS T2 ON T1.superhero_name = T2.superhero_name
WHERE T2.publisher_name = ’Marvel’;
For the above hybrid query example, the user wants to know the full name and superhero name for all heroes from the Marvel universe; However, the publisher information is missing from the database. Assume the publisher information has been extracted and materialized in a (virtual) $llm$ table; then, this hybrid query can answer the user’s question.
Formally, a hybrid query $q_{hybrid}$ answers a beyond-database question $Q$ over database $D$ and the (virtual) table generated from an LLM, $f(T|\theta)$, where $\theta$ represents the parameters of the model. Hence the hybrid query task can be formulated as: $q_{hybrid}(D,f(T|\theta))$.
@@ 2.3. Related Work
Answering beyond-database questions has been investigated in a variety of research efforts. CrowdDB (Franklin et al., 2011) introduces a crowdsourced query processing system that addresses the closed-world assumption in traditional query processing. However, the cost of incorporating human input can be significant, impacting both time and resources. Inspired by the capability of LLMs, researchers have investigated whether LLMs can be used for data cleaning tasks such as data imputation where LLMs repair dirty or missing values in data entries (Narayan et al., 2022; Chen et al., 2023). While this is closely related to hybrid queries, hybrid querying over relational databases and LLMs presents its own set of challenges in combining structured and unstructured data sources, ensuring the correctness of generated data, and materializing the data for future uses. Furthermore, two recent works (Urban et al., 2023; Saeed et al., 2023) laid out the vision for augmenting relational databases with data generated from LLMs. However, due to the absence of an evaluation benchmark, both studies are limited to preliminary case studies. For instance, Galois (Saeed et al., 2023) executed 46 SQL queries (drawn from the Spider benchmark (Yu et al., 2018)) solely on LLMs, without involving any relational databases. Also, they manually verified the generation results, comparing the output table statistics (e.g., cardinality) and verifying content accuracy with the ground truth. Our proposed SWAN benchmark is built on top of databases collected in the Bird benchmark (Li et al., 2024), which has 270x (on average) more rows compared to the databases in the Spider benchmark (Yu et al., 2018). Moreover, in addition to the content accuracy, we also compare the execution accuracy among hybrid queries (see more explanations in Section 5)
@@ 3.1. Databases in SWAN
We constructed the SWAN benchmark111see https://github.com/ZhaoFuheng/SWAN/ based on the Bird benchmark (Li et al., 2024). In the Bird benchmark, there are eleven diverse database domains. However, we have identified that many of these databases are overly narrow, and the questions are too specific to be answered effectively by a general intelligence model. For instance, the financial database includes detailed tables on bank accounts, credit card information, loans, and trading transactions. This level of specificity are out of the scope of a general AI’s capabilities. As a result, we selected four diverse databases: European Football, Formula One, California Schools, and Superhero. These databases cover a broad range of topics, from sports statistics and history to educational trends and fictional characters.
@@ 3.2. Schema Curation
The main challenges in evaluating hybrid queries are: i) generating beyond-database questions, and ii) ensuring the availability of ground truth answers for these questions. Fortunately, the Bird benchmark provides valuable assets: natural language questions, SQL queries, and the databases. Leveraging these resources, we can modify the databases to create questions that the databases can not answer based on the database content. For the four selected databases, we removed specific columns or entire tables to generate beyond-database questions. For example, in the Superhero database, the $superhero$ table contains a $publisher\_id$ field, which is a foreign key used to identify the $publisher\_name$ in the $publisher$ table. By dropping the $publisher\_id$ column, all questions related to finding the name of the publisher become unanswerable based on the newly curated database. After the schema curation, the statistics of the selected databases are shown in Table 1.
@@
Formula One 13 39560.5 12
California Schools 3 9980.3 12
Superhero 10 1061.4 11
@@ 3.3. Free Form Response and Value Selection
In SWAN, the challenges for LLMs to generate factual data can be broken down into two categories: i) free form response and ii) value selection. The free form response requires LLMs to generate data when some context is provided. For instance, in the California Schools database, it originally contained both the school name and the school url. We removed the school url column, and as a result, we expect LLMs to generate short-form urls for the schools. Often, the school url is closely related to the school name and often ends with edu. The value selection involves choosing data values from a predefined list. For example, when we removed the $publisher\_id$ field from the Superhero database, we retained the $publisher$ table, which contains the names of all publishers for the superheroes in the database. Consequently, the list of all publisher names can be provided to the LLMs, allowing them to select the appropriate publisher for each superhero.
@@ 3.4. Keys for Tables from LLMs
In relational databases, a foreign key column is often represented as an integer linked to the primary key column in another table. However, intergers do not provide any meaningful insights for LLMs to generate useful data values. According to SQL standards, a foreign key must reference a unique key in the foreign table. Therefore, we have curated the databases to include meaningful foreign keys for the data generated by LLMs. For example, in the Superhero database, we assume the combination of $superhero\_name$ and $full\_name$ of a superhero serves as the key to finding the publisher information. Also, we have ensured that there are no duplicate pairs of ($superhero\_name$, $full\_name$) in the table.
@@ 3.5. Beyond-Database Questions
For each database in SWAN, we provide 30 beyond-database questions, resulting in a total of 120 questions across all databases222see https://github.com/ZhaoFuheng/SWAN/. For each question, we also supply a hybrid SQL query that joins the tables in the relational database with the tables generated by LLMs. It is important to note that these hybrid queries assume the values generated by LLMs are materialized as tables.
4. HQDL
In this section, we discuss our proposed Hybrid Query Database LLM (HQDL), a preliminary solution for solving beyond-database questions with hybrid queries.
4.1. Data Generation
For each database, SWAN provides a list of missing columns and tables that need to be generated by LLMs. In addition, SWAN specifies the keys consisting of the minimal number of attributes that represent the primary-key/foreign-key (PK-FK) relationships between the existing tables in the relational database and the tables generated by LLMs. This ensures that the necessary keys can be used as input for the LLMs, enabling them to accurately generate the missing data entries. A zero-shot prompt for generating the missing data values for the Superhero database is provided below:
@@ 4.1. Data Generation
HQDL needs to instruct the LLMs to fill in the missing values in the target data entry. HQDL also adopts the widely accepted ’No Exaplanation’ rule introduced by OpenAI (OpenAI, 2024) and the ’No Exaplanation’ consistently improves the quality of generated answer for semantic parsing (Gao et al., 2023). Furthermore, HQDL provides value lists, such as publishers and colors, for LLMs to select from. Since only the id fields are removed (e.g., $publisher\_id$), HQDL can directly retrieve all these predefined data values. The goal is to avoid ambiguous data values such as ’Marvel’ v.s. ’Marvel Comics’ in which both represent the same publisher but pose challenges for automatic evaluation.
@@
Instructions.
Value lists of publishers, colors, power names, etc.
/*One example is provided before the target data entry*/
@@
As shown above, one example data entry and the corresponding answer are provided about the ’3-D Man’. In the evaluation section (Section 5), we will show that few-shot demonstrations significantly improve the quality of the generated data entries.”
4.2. Data Extraction
After collecting all data entries generated by the LLMs, HQDL materializes these entries into tables. HQDL uses the Python csv module’s reader to process these entries, converting them into a structured format, and inserting them into new tables in the underlying SQLite database. Moreover, in SWAN, there are both one-to-one and one-to-many relationships. When one-to-many relationships occur, HQDL condenses the tuples in the ”many” side of the relationship into a long text. For example, each superhero may be associated with many powers. HQDL would condense all the powers into a long string separated by commas (e.g., ”Agility, Super Strength,Super Speed”).
4.3. Limitations of HQDL
While the preliminary solution, HQDL, can be used to solve the challenges presented in our proposed SWAN benchmark, there are a few limitations. First, HQDL does not provide more context than the necessary keys and the predefined value lists. There are other attributes inside the relational database that may be relevant and it remains an open question on how to select the best context. Second, the prompts and static examples used in HQDL are hand-crafted. It would be more convenient for users if the data system may automatically generate prompts and examples based on the specific context and query requirements. Third, HQDL requires LLMs to generate all deleted data and also materialize these data as tables. Since the goal is to answer beyond-database questions, generating all deleted data may not be necessary. Lastly, HQDL separates the pipeline into small steps: i) generating all deleted data with static prompts; ii), materializing this data into tables; and iii), executing the hybrid queries to obtain answers for the beyond-database questions. We observe that the industry has started integrating LLM calls directly into SQL syntax. For instance, Google BigQuery provides syntax to directly query LLMs using SQL (Kashyap, 2023). Hence, finding all hero names from the Marvel universe within the database can be rewritten as follows:

### v2

@@ Abstract.
Database queries traditionally operate under the closed-world assumption, providing no answers to questions that require information beyond the data stored in the database. Hybrid querying using SQL offers an alternative by integrating relational databases with large language models (LLMs) to answer beyond-database questions. In this paper, we present the first cross-domain benchmark, SWAN, containing 120 beyond-database questions over four real-world databases. To leverage state-of-the-art language models in addressing these complex questions in SWAN, we present two solutions: one based on schema expansion and the other based on user defined functions. We also discuss optimization opportunities and potential future directions. Our evaluation demonstrates that using GPT-4 Turbo with few-shot prompts, one can achieves up to 40.0% in execution accuracy and 48.2% in data factuality. These results highlights both the potential and challenges for hybrid querying. We believe that our work will inspire further research in creating more efficient and accurate data systems that seamlessly integrate relational databases and large language models to address beyond-database questions.
@@ 1. Introduction
The Relational model and SQL have achieved widespread acceptance and usage in data management systems. In particular, SQL continues to evolve by adding new syntax and features, enhancing its capabilities to meet the growing demands of modern data systems (Stonebraker and Pavlo, 2024). It is well known that database queries are evaluated under a closed domain assumption, meaning that the queries address aspects of the real world based solely on the data stored in the relational database management system (Reiter, 1988). While the direct approach is to say NO to beyond-database questions, researchers in our community have also investigated alternatives such as providing answers based on incomplete data with heuristic algorithms (Levy, 1996) and crowdsourcing (Franklin et al., 2011). In this paper, we explore the use of large language models to address these questions.
In the natural language processing community, open-domain question answering has been extensively studied, often encompassing a broader range of inquiries. This long-standing task aims to provide factual answers to natural language questions by drawing from large, unstructured collections of texts and documents, such as Wikipedia. By pre-training on large corpus of knowledge, large language models (LLMs) have demonstrated significant potential in providing world knowledge and performing complex reasoning (Hendrycks et al., 2020; Zhao et al., 2023). In this paper, we are specifically interested in beyond-database questions which have partial information grounded in the relational database. Unlike open-domain questions, beyond-database questions require integrating and reasoning with structured data from relational databases as well as drawing from large, unstructured collections of texts.
@@
A motivating example: Consider a simple database with a single table containing superhero information, as shown in Figure 1. The schema for the database is: superhero(hero_name, full_name). Suppose the user wants to list all the hero names from the Marvel Universe within the database. This would be considered as a beyond-database question, since the database contains relevant or partial information to the request but cannot directly provide the answer (i.e., which heroes are Marvel universe). On the other hand, large language models such as ChatGPT can be used to identify the publisher for each superhero characters. Assume we treat LLMs as a table containing the hero_name and the publisher. Then, a SQL query like: ”SELECT hero_name, full_name FROM LLM JOIN superhero ON LLM.hero_name = superhero.hero_name WHERE llm.publisher = ’Marvel’;” Hence, hybrid querying by integrating relational databases and large language models offers a powerful approach for addressing beyond-database questions.
In this work, we propose SWAN, Solving beyond-database queries With generative AI aNd relational databases, the first hybrid query benchmark. SWAN comprises 120 beyond-database questions and spans four big databases, covering four diverse domains. The original databases and questions are from the recent Bird benchmark(Li et al., 2024a), which is a benchmark for evaluating natural language to SQL translation. The data in these databases are collected from open-source relational databases in platforms such as Kaggle. Also, the proposed SWAN benchmark challenges large language models to both select values from a given list (e.g., choose publisher name from a list of predefined publishers) and generate free-form outputs (e.g., determine the city based on a street address). In addition to proposing the SWAN benchmark, we also introduce HQDL, a preliminary solution for answering beyond-database questions. We evaluate HQDL over the state-of-the-art large language models such as ChatGPT (gpt-3.5-turbo) (Ouyang et al., 2022) and the GTP-4 turbo (Open, 2023; Achiam et al., 2023). Our experimental results indicate significant challenges for current state-of-the-art models in generating factual data and accurately answering beyond-database questions with hybrid queries. With in-context learning (ICL), GPT-4 Turbo achieves 40.0% execution accuracy and 48.2% of its generated data is factually correct. Proposing a novel benchmark for beyond-database queries will focus attention and drive more research in this critical area. It will also encourage researchers to develop innovative solutions that are founded on solid well understood foundations. The evaluation results based on HQDL further highlight the need for more advancements in improving the accuracy and reliability of answering beyond-database questions with hybrid queries.
@@ 2. Background and Problem Statement
In this section, we provide some basic background of large language models (LLMs) and introduce related work from both the database and the natural language communities for answering beyond-database questions.
2.1. Large Language Models Basics
@@
Large language models are frequently used to retrieve domain-specific knowledge and to simplify the information retrieval process by providing direct natural language answers. However, these models may exhibit hallucinations and factuality errors. Consequently, recently many researchers have focused on enhancing the capabilities of LLMs to provide factual information (Wang et al., 2023a). In-context learning (ICL) is a widely adopted strategy to enhance the factual accuracy of generated content. ICL enables a language model to learn from example demonstrations within its context to improve performance (Brown et al., 2020).
@@ 2.2. Related Work
Answering beyond-database questions has been investigated in a variety of research efforts. CrowdDB (Franklin et al., 2011), Qurk (Marcus et al., 2011), Deco (Park et al., 2012), and hQuery (Parameswaran and Polyzotis, 2011) introduce crowdsourced query processing systems that address the closed-world assumption in traditional query processing. However, the cost of incorporating human input can be significant, impacting both time and resources. Inspired by the capabilities of LLMs, researchers have investigated whether LLMs can be used for declarative prompting (Parameswaran et al., 2023) and data cleaning tasks such as data imputation where LLMs repair dirty or missing values in data entries (Narayan et al., 2022; Chen et al., 2023). While this is closely related to hybrid queries, hybrid querying over relational databases and LLMs presents its own set of challenges in combining structured and unstructured data sources, ensuring the correctness of generated data, and materializing the data for future uses. Furthermore, two recent studies (Urban et al., 2023; Saeed et al., 2023) laid out the vision for augmenting relational databases with data generated from LLMs. However, due to the absence of an evaluation benchmark, both studies are limited to preliminary case studies. For instance, Galois (Saeed et al., 2023) executed 46 SQL queries (drawn from the Spider benchmark (Yu et al., 2018)) solely on LLMs, without involving any relational databases. Also, they manually verified the generation results, comparing the output table statistics (e.g., cardinality) and verifying content accuracy with the ground truth. Our proposed SWAN benchmark is built on top of databases collected in the Bird benchmark (Li et al., 2024a), which has 270x (on average) more rows compared to the databases in the Spider benchmark (Yu et al., 2018). Moreover, in addition to content accuracy, we also compare the execution accuracy among hybrid queries (see more explanations in Section 5)
@@ 3.1. Databases in SWAN
We constructed the SWAN benchmark111see https://github.com/ZhaoFuheng/SWAN/ based on the Bird benchmark (Li et al., 2024a). In the Bird benchmark, there are eleven diverse database domains. However, we have identified that many of these databases are overly narrow, and the questions are too specific to be answered effectively by a general intelligence model. For instance, the financial database includes detailed tables on bank accounts, credit card information, loans, and trading transactions. This level of specificity are out of the scope of a general AI’s capabilities. As a result, we selected four diverse databases: European Football, Formula One, California Schools, and Superhero. These databases cover a broad range of topics, from sports statistics and history to educational trends and fictional characters.
@@ 3.2. Schema Curation
The main challenges in evaluating hybrid queries are: i) generating beyond-database questions, and ii) ensuring the availability of ground truth answers for these questions. Fortunately, the Bird benchmark provides valuable assets: natural language questions, SQL queries, and the databases. Leveraging these resources, we can modify the databases to create questions that the databases can not answer based on the database content. For the four selected databases, we removed specific columns or entire tables to generate beyond-database questions. For example, in the Superhero database, the $superhero$ table contains a $publisher\_id$ field, which is a foreign key used to identify the $publisher\_name$ in the $publisher$ table. By dropping the $publisher\_id$ column, all questions related to finding the name of the publisher become unanswerable based on the newly curated database. While the entire publisher table can be directly dropped, we kept the distinct values of $publisher\_name$ to assist LLMs in correctly formatting the output related to publishers (see more explanation in the next section). After schema curation, the statistics of the selected databases are shown in Table 1.
@@
Formula One 13 39561 12
California Schools 3 9980 12
Superhero 10 1061 11
@@ 3.3. Free Form Response and Value Selection
In SWAN, the challenges for LLMs to generate factual data can be broken down into two categories: i) free form response and ii) value selection. The free form response requires LLMs to generate data when some context is provided. For instance, in the California Schools database, the tables originally contained both the school name and the school url. We removed the school url column, and as a result, we expect LLMs to generate short-form urls for the schools. Often, the school url is closely related to the school name and often ends with edu. Value selection involves choosing data values from a predefined list (e.g., a list of unique publisher names). For instance, after we removed the $publisher\_id$ field from the $superhero$ table and the entire $publisher$ table from the Superhero database, we retained the unique values of $publisher\_names$, which contains the names of all publishers for the superheroes in the database. Consequently, the list of all publisher names can be provided to the LLMs, allowing them to select the appropriate publisher for each superhero.
@@ 3.4. Keys for Tables from LLMs
In relational databases, a foreign key column is often represented as an integer linked to the primary key column in another table. However, integers do not provide any meaningful insights for LLMs to generate useful data values. According to SQL standards, a foreign key must reference a unique key in the foreign table. Therefore, we have curated the databases to include meaningful foreign keys for the data generated by LLMs. For example, in the Superhero database, we assume the combination of $superhero\_name$ and $full\_name$ of a superhero serves as the key to finding the publisher information. Also, we have ensured that there are no duplicate pairs of ($superhero\_name$, $full\_name$) in the table. Our approach of designing meaningful keys for LLMs to generate data aligns with the data model in crowd-sourcing systems. For example, Deco’s Fetch/Resolution rules (Park et al., 2012) use meaningful keys as input and ask crowd-source workers to generate a group of attributes based on the given keys.
@@ 3.5. Beyond-Database Questions
For each database in SWAN, we provide 30 beyond-database questions, resulting in a total of 120 questions across all databases. For each question, we also supply i) a hybrid SQL query that joins the tables in the relational database with the tables generated by LLMs (assume the values generated by LLMs are materialized as tables), ii) a hybrid SQL query that directly invokes LLM calls based on BlendSQL (Glenn et al., 2024) functions, and iii) the corresponding gold SQL query from Bird, such that the expected answer is the execution results of the gold SQL query on the original Bird databases. Notably, these hybrid queries are manually crafted and fully executable. They are provided to assess the current capabilities of combining LLMs with databases to answer beyond-database questions. Automating the translation of beyond-database questions into hybrid queries is left as future work.
4. Answering Beyond-Database Questions
In this section, we discuss two different approaches to answer beyond-database questions in SWAN. One is based on schema expansion and the other is based on SQL user defined functions. At the end of this section, we discuss the promising optimization opportunities of these solutions and outline potential directions for future research.
4.1. HQDL
First, we introduce Hybrid Query Database and LLM (HQDL), a preliminary solution for solving beyond-database questions based on schema expansion. Given a beyond-database $NL$ question, one can expand the database schema by including new columns or new tables such that $NL$ question becomes answerable based on the new schema. Then, LLMs can be used to fill in all the missing data entries after schema expansion. Based on the newly updated schema, one can write a regular SQL query to answer question $NL$ directly.
4.1.1. Data Generation
For each database, SWAN provides a list of missing columns and tables that need to be generated by LLMs. SWAN has also provided the keys consisting of the minimal number of attributes that represent the primary-key/foreign-key (PK-FK) relationships between the existing tables in the relational database and the tables generated by LLMs. This ensures that the necessary keys can be used as input for the LLMs, enabling them to accurately generate the missing data entries. We would also like to note that these information can be helpful, though its use is optional. In HQDL, we choose to directly use these metadata information. An example of a zero-shot prompt used to generate missing data values in the Super Hero database is provided below. This prompt instructs LLMs to infer and fill in missing data entries by supplying guidance, column names, and example values for certain columns (e.g., publishers and colors).
@@ 4.1.1. Data Generation
HQDL needs to instruct the LLMs to fill in the missing values in the target data entry. HQDL also adopts the widely accepted ’No Explanation’ rule introduced by OpenAI (OpenAI, 2024), which consistently improves the quality of generated answers for semantic parsing (Gao et al., 2023). Furthermore, HQDL provides value lists, such as publishers and colors, for LLMs to select from. Since only the id fields are removed (e.g., $publisher\_id$), HQDL can directly retrieve all these predefined data values. The goal is to avoid ambiguous data values such as ’Marvel’ v.s. ’Marvel Comics’ in which both values represent the same publisher but pose challenges for automatic evaluation.
@@
Prefix (instructions and value lists)
/*An example is provided before the target data entry*/
@@
As shown above, an example data entry and the corresponding answer are provided to the LLM for the constructed record corresponding to ’3-D Man’ and Charles Chandler. In the evaluation section (Section 5), we will show that few-shot demonstrations significantly improve the quality of the generated data entries.”
Data Extraction. After collecting all data entries generated by the LLMs, HQDL materializes these entries into tables. HQDL uses the Python csv module’s reader to process these entries, converting them into a structured format, and inserting them into new tables in the underlying SQLite database. Moreover, in SWAN, there are both one-to-one and one-to-many relationships. When one-to-many relationships occur, HQDL condenses the tuples in the ”many” side of the relationship into a long text. For example, each superhero may be associated with many powers. HQDL would condense all the powers into a long string separated by commas (e.g., ”Agility, Super Strength, Super Speed”).
4.2. Hybrid Query UDFs
We observe that industry has started integrating LLM calls directly into SQL syntax through user defined functions, such as DucksDB (Raasveldt and Mühleisen, 2019) and Google BigQuery (Kashyap, 2023). For instance, finding all hero names from the Marvel universe within the database can be rewritten as follows in Google BigQuery: